OpenStudy (anonymous):

Find x3 − 2y2 − 3x3 + z4 if x = 3, y = 5, and z = −3.

OpenStudy (anonymous):

Find x3 − 2y^2 − 3x^3 + z^4 if x = 3, y = 5, and z = −3.

OpenStudy (anonymous):

3^3 - 2(5)^2 - 3(3)^3 + (-3)^4

OpenStudy (astrophysics):

Yes so now evaluate

OpenStudy (astrophysics):

\[3^3 = 3 \times 3 \times 3\] \[2(5)^2 = 2 (5 \times 5)\] \[-3(3)^3 = -3(3 \times 3 \times 3)\] \[(-3)^4 = (-3 \times -3 \times - 3 \times - 3)\] see the pattern here?